OpenStudy (anonymous):

write equation for parabola focus (2,1) opens right, contains (8,-7)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

k i find the distance between F and point given which is 10 then what?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

ok here's a good picture that describes what a parabola truly is http://www.mathwords.com/f/focus_parabola.htm

jimthompson5910 (jim_thompson5910):

it's basically a set of points that if you pick any point on the parabola P, then the distance from P to F (focus) will be the same as the distance from P to the directrix

jimthompson5910 (jim_thompson5910):

you found the distance from the focus to the point to be 10 units

jimthompson5910 (jim_thompson5910):

so this point is 10 units from the directrix

jimthompson5910 (jim_thompson5910):

so where is the directrix? what's the equation?

OpenStudy (anonymous):

hold on

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

i will move left 10 points?

OpenStudy (anonymous):

confused

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

about what

OpenStudy (anonymous):

can we work throught this question?

jimthompson5910 (jim_thompson5910):

if x = 8 and you go 10 units to the left, then you end up at x = ??

OpenStudy (anonymous):

-2?

jimthompson5910 (jim_thompson5910):

so x = -2 is the directrix

jimthompson5910 (jim_thompson5910):

the vertex will be the point (0,1) since that's the midpoint of the segment that connects the focus and it's perpendicular to the directrix

OpenStudy (anonymous):

hmm can u plz plz plz draw that?

jimthompson5910 (jim_thompson5910):

sry it's hard to draw a parabola on here

OpenStudy (anonymous):

it is okay k next step

jimthompson5910 (jim_thompson5910):

you know the vertex is (0,1), so h = 0, k = 1

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the focal distance is p = 2 since the vertex is 2 units way from the directrix

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@jim_thompson5910 that's it?

jimthompson5910 (jim_thompson5910):

no you would use the formulas given on that page i sent you

jimthompson5910 (jim_thompson5910):

to find the equation

OpenStudy (anonymous):

okay i will do that!

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (anonymous):

(y-1)^2=8(x-0)

jimthompson5910 (jim_thompson5910):

perfect, you can also write that as (y-1)^2 = 8x

OpenStudy (anonymous):

okay thanks!

jimthompson5910 (jim_thompson5910):

yw